OpenStudy (anonymous):

2.4=3 (m + 4) show work

OpenStudy (anonymous):

2.4 = 3m + 12 2.4/3 = 3m/3 + 12/3 .8 = m + 4 .8 - 4 = m + 4 - 4 -3.2 = m

OpenStudy (anonymous):

thats what i got